OpenStudy (bohotness):

A children's game involves tossing a bean bag onto a lettered board. The board is divided into six sections of equal size. The sections are labeled A through F. If a bean bag is randomly tossed onto the board, what is the probability that it will land on B or C?

OpenStudy (igreen):

6 sections. Probability it will land on B or C.

OpenStudy (bohotness):

https://media.studyisland.com/pics/162727a.png

OpenStudy (igreen):

B or C is how many sections?

OpenStudy (igreen):

2, right? So the probability is 2/6. Simplify 2/6

OpenStudy (bohotness):

yes

OpenStudy (bohotness):

1. A children's game involves tossing a bean bag onto a lettered board. The board is divided into six sections of equal size. The sections are labeled A through F. If a bean bag is randomly tossed onto the board, what is the probability that it will land on B or C? https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{1}{6} https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{2}{3} https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{1}{36} https://media.studyisland.com/cgi-bin/mimetex.cgi? \frac{1}{3}

OpenStudy (bohotness):

@iGreen